OpenStudy (anonymous):

(y-8)/(y-9)=(1)/(y-9) y=?

OpenStudy (jlastino):

9 Just multiply both sides by (y-9) to cancel the denominators y-8=1 y=1+8 y=9

jimthompson5910 (jim_thompson5910):

answer is not 9 since y can't equal 9 (to avoid division by zero)